# On relative tractor bundles

### v1

@@ 3.1. Restrictions of the adjoint tractor connection
A crucial feature of tractor bundles is that they carry canonical linear connections known as tractor connections. There is a uniform description of these tractor connections, based on the so-called fundamental derivative. For a geometry $(p:\mathcal{G}\to M,\omega)$ of type $(G,Q)$ consider any representation $\mathbb{W}$ of $Q$ and the corresponding natural vector bundle $\mathcal{W}M:=\mathcal{G}\times_{Q}\mathbb{W}$. Then smooth sections of $\mathcal{W}M$ are in bijective correspondence with smooth functions $\mathcal{G}\to\mathbb{W}$ which are $Q$-equivariant in the sense that $f(u\cdot g)=g^{-1}\cdot f(u)$ for any $g\in Q$, where in the left hand side we use the principal right action of $Q$ on $\mathcal{G}$ and in the right hand side the given representation on $Q$. In this way, sections of the adjoint tractor bundle $\mathcal{A}M$ get identified with $Q$-equivariant smooth functions $f:\mathcal{G}\to\mathfrak{g}$. But via the Cartan connection $\omega$, smooth functions $\mathcal{G}\to\mathfrak{g}$ are in bijective correspondence with vector fields on $\mathcal{G}$, and the equivariancy condition on $f$ is equivalent to the corresponding vector field $\xi\in\mathfrak{X}(\mathcal{G})$ being invariant under the principal right action of $Q$, i.e. $(r^{g})^{*}\xi=\xi$ for any $g\in Q$.
@@
Differentiating an equivariant function with respect to an invariant vector field produces an equivariant function, so we can view this as defining an operator $D:\Gamma(\mathcal{A}M)\times\Gamma(\mathcal{W}M)\to\Gamma(\mathcal{W}M)$ called the fundamental derivative, see Section 1.5.8 of [book]. This has strong invariance properties and to emphasize the analogy to a covariant derivative, it is usually written as $(s,\sigma)\mapsto D_{s}\sigma$.
@@
While things are not formulated in this way in [Rel-BGG2] (see the proof of Theorem 5 for details), relative tractor connections can be constructed in a very similar fashion. Suppose we have a bundle $\mathcal{V}M$ that is induced by a representation of $P$, where $Q\subset P\subset G$ as in §2.2. Then for any $g\in Q$, the subspace $\mathfrak{p}\subset\mathfrak{g}$ is invariant under $\operatorname{Ad}(g):\mathfrak{g}\to\mathfrak{g}$, so it gives rise to a smooth subbundle in $\mathcal{A}M$ that is denoted by $\mathcal{A}^{\mathfrak{p}}M$ in [Rel-BGG2]. The infinitesimal representation again induces a bundle map $\bullet:\mathcal{A}^{\mathfrak{p}}M\times\mathcal{V}M\to\mathcal{V}M$. As before, we conclude that the operator $\Gamma(\mathcal{A}^{\mathfrak{p}}M)\times\Gamma(\mathcal{V}M)\to\Gamma(\mathcal{V}M)$ defined by $(s,\sigma)\mapsto D_{s}\sigma+s\bullet\sigma$ depends only on $\Pi(s)$, but now $\Pi(s)$ is a section of $\mathcal{G}\times_{Q}(\mathfrak{p}/\mathfrak{q})=T_{\rho}M$. Hence we don’t obtain a full linear connection but only a partial connection $\nabla^{\rho,\mathcal{V}}:\Gamma(T_{\rho}M)\times\Gamma(\mathcal{V}M)\to\Gamma(\mathcal{V}M)$, which still is linear over smooth functions in the first argument and satisfies the Leibniz rule in the second argument.
In the special case of the relative tractor bundles $\mathcal{V}_{i^{\prime}}M$ constructed in §2.3 we prove that the relative tractor connection is induced by the adjoint tractor connection.
Theorem 5.
In the setting of §2.3, consider the subbundles $\mathcal{A}^{(i^{\prime},*)}M$. Then for $s\in\Gamma(\mathcal{A}^{(0,*)}M)$ and $t\in\Gamma(\mathcal{A}^{(i^{\prime},*)}M)$ with $i^{\prime}\leq 0$ we get $\nabla^{\mathcal{A}}_{\Pi(s)}t\in\Gamma(\mathcal{A}^{(i^{\prime},*)})M$. Hence the adjoint tractor connection restricts to an operation $\Gamma(T_{\rho}M)\times\Gamma(T^{i^{\prime}}_{P}M)\to\Gamma(T^{i^{\prime}}_{P}M)$ for any $i^{\prime}<0$.
Thus there is an induced partial connection on $\Gamma(\mathcal{V}_{i^{\prime}}M)$ and this coincides with the relative tractor connection $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}$ as obtained in Section 4.3 of [Rel-BGG2].
@@
Consider the defining formula $\nabla^{\mathcal{A}}_{\Pi(s)}t=D_{s}t+\{s,t\}$ for $s\in\Gamma(\mathcal{A}^{(0,*)}M)$ and $t\in\Gamma(\mathcal{A}^{(i^{\prime},*)}M)$. By Proposition 1.5.8 of [book], the fundamental derivative preserves natural subbundles, so the first summand in the right hand side is a section of $\mathcal{A}^{(i^{\prime},*)}M$. On the other hand, Proposition 4 together with $\mathfrak{p}=\mathfrak{g}^{(0,*)}$ shows that $[\mathfrak{g}^{(0,*)},\mathfrak{g}^{(i^{\prime},*)}]\subset\mathfrak{g}^{(i^{\prime},*)}$, so the second summand of the right hand side of the defining formula is a section of $\mathcal{A}^{(i^{\prime},*)}M$, too.
@@
Fixing $i^{\prime}<0$, we have thus obtained a partial connection on the bundle $\mathcal{A}^{(i^{\prime},*)}M$ for which the subbundle $\mathcal{A}^{(i^{\prime}+1,*)}M$ is parallel, so it descends to a well defined partial connection on the quotient bundle $\mathcal{V}_{i^{\prime}}M=\mathcal{A}^{(i^{\prime},*)}M/\mathcal{A}^{(i^{\prime}+1,*)}M$. Thus it remains to show that this coincides with the relative tractor connection. The construction of the latter in [Rel-BGG2] was based on a variant of the fundamental derivative for relative natural bundles: One defines the relative adjoint tractor bundle $\mathcal{A}_{\rho}M=\mathcal{G}\times_{Q}(\mathfrak{p}/\mathfrak{p}_{+})$, so in our notation, this is just $\mathcal{A}^{(0,*)}M/\mathcal{A}^{(1,*)}M$. Given a representation $\mathbb{W}$ of $Q$ on which $P_{+}$ acts trivially, on can first restrict the fundamental derivative $D_{s}\sigma$ to $s\in\Gamma(\mathcal{A}^{(0,*)}M)$ and then observe that $D_{s}\sigma=0$ if $s\in\Gamma(\mathcal{A}^{(1,*)}M)$, to obtain a well defined operator $D^{\rho}:\Gamma(\mathcal{A}_{\rho}M)\times\Gamma(\mathcal{W}M)\to\Gamma(\mathcal{W}M)$. If $\mathcal{V}M$ is a relative tractor bundle, then $\mathbb{V}$ is the restriction of a representation of $P$, so the infinitesimal representation descends to an operation $\bullet:\mathfrak{p}/\mathfrak{p}_{+}\times\mathbb{V}\to\mathbb{V}$. Similarly as above, the relative tractor connection is then characterized by $\nabla^{\rho,\mathcal{V}}_{\Pi(s)}\sigma=D^{\rho}_{s}\sigma+s\bullet\sigma$, where now $\Pi$ is induced by the canonical projection $\mathfrak{p}/\mathfrak{p}_{+}\to\mathfrak{p}/\mathfrak{q}$.
@@
But from this description the claim follows easily: Take sections $\sigma\in\Gamma(\mathcal{V}_{i^{\prime}}M)$ and $\xi\in T_{\rho}M$. For our construction, we choose representative sections $t\in\Gamma(\mathcal{A}^{(i^{\prime},*)}M)$ for $\sigma$ and $s\in\Gamma(\mathcal{A}^{(0,*)}M)$ for $\xi$ and project $D_{s}t+\{s,t\}\in\Gamma(\mathcal{A}^{(i^{\prime},*)}M)$ to the quotient bundle $\mathcal{V}_{i^{\prime}}M$. Denoting by $\underline{s}$ the projection of $s$ to $\mathcal{A}_{\rho}M$ the fact that $\mathfrak{p}_{+}$ acts trivially on $\mathfrak{g}^{(i^{\prime},*)}/\mathfrak{g}^{(i^{\prime}+1,*)}$ implies that the projection of $D_{s}t$ to $\Gamma(\mathcal{V}_{i^{\prime}}M)$ depends only on $\underline{s}$ and then by construction it coincides with $D^{\rho}_{\underline{s}}\sigma$. Similarly, since $[\mathfrak{g}^{(1,*)},\mathfrak{g}^{(i^{\prime},*)}]\subset\mathfrak{g}^{(i^{\prime}+1,*)}$ we conclude that the projection of $\{s,t\}$ to $\Gamma(\mathcal{V}_{i^{\prime}}M)$ depends only on $\underline{s}$ and coincides with $\underline{s}\bullet\sigma$. ∎
Remark 6.
@@
The description of a relative tractor connection as induced by the adjoint tractor connection leads to explicit descriptions in terms of the structure underlying a parabolic geometry of type $(G,Q)$. This works via the concept of Weyl structures which were introduced in [Weyl], see also Chapter 5 of [book]. Any parabolic geometry carries a family of Weyl structures which form an affine space modelled on the space $\Omega^{1}(M)$ of one-forms on the manifold $M$. Any Weyl structure provides a linear connection on any associated vector bundle called the Weyl connection. Once Weyl connections are understood there is a general theory providing explicit formulae for tractor connections, see Section 5.1 of [book].
3.2. The relation to the Lie bracket
@@
As we have noted above, sections of the adjoint tractor bundle $\mathcal{A}M$ can be identified with $Q$-invariant vector fields on the total space $\mathcal{G}$ of the Cartan bundle. Now the subspace of $Q$-invariant vector fields is closed under the Lie bracket of vector fields, whence we obtain an induced bilinear operation on $\Gamma(\mathcal{A}M)$, which we denote by the usual symbol [,][\ ,\ ][ , ]. As we have noted in §3.1, for $s\in\Gamma(\mathcal{A}M)$, $\Pi(s)\in\mathfrak{X}(M)$ is the projection of the $Q$-invariant vector field corresponding to $s$. Hence standard properties of the Lie bracket imply that for $s,t\in\Gamma(\mathcal{A}M)$, we get $\Pi([s,t])=[\Pi(s),\Pi(t)]$ with the Lie bracket of vector fields in the right hand side.
@@
It is rather easy to express the Lie bracket of sections of $\mathcal{A}M$ in terms of the operations introduced above and of the curvature $\kappa$ of the Cartan connection $\omega$, see Corollary 1.5.8 of [book]: For $s,t\in\Gamma(\mathcal{A}M)$, one has
(3.1) $[s,t]=D_{s}t-D_{t}s-\kappa(\Pi(s),\Pi(t))+\{s,t\}.$
@@
Using this and Theorem 5, we can now give an alternative description of the relative tractor connections on the relative tractor bundles $\mathcal{V}_{i^{\prime}}M$. Recall from §2.3 that we have a subbundles $T^{i^{\prime}}_{P}M\subset T_{\rho}M\subset TM$ such that $\mathcal{V}_{-1}M=T^{-1}_{P}M/T_{\rho}M$ and $\mathcal{V}_{i^{\prime}}M=T^{i^{\prime}}_{P}M/T^{i^{\prime}+1}_{P}M$ for $i^{\prime}<-1$. Hence for a smooth section $\sigma$ of $\mathcal{V}_{i^{\prime}}M$ one can always chose a representative vector field in $\Gamma(T^{i^{\prime}}_{P}M)$ that projects onto $\sigma$.
Theorem 7.
@@
In the setting of §2.3 consider vector fields $\xi\in\Gamma(T_{\rho}M)$, $\eta\in\Gamma(T^{i^{\prime}}_{P}M)$ and a section $\sigma\in\Gamma(\mathcal{V}_{i^{\prime}}M)$. Let $\tau\in\Omega^{2}(M,TM)$ be the torsion of the Cartan connection $\omega$.
(1) The vector field $[\xi,\eta]+\tau(\xi,\eta)$ is a section of the subbundle $T^{i^{\prime}}_{P}M\subset TM$.
@@
(2) If $\eta$ is a representative for $\sigma$, then the projection of $[\xi,\eta]+\tau(\xi,\eta)$ to $\Gamma(\mathcal{V}_{i^{\prime}}M)$ depends only on $\sigma$ (and not on the choice of $\eta$) and coincides with $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}_{\xi}\sigma$.
@@
We can choose smooth representatives $s\in\Gamma(\mathcal{A}^{(0,*)}M)$ for $\xi$ and $t\in\Gamma(\mathcal{A}^{(i^{\prime},*)}M)$ for $\eta$, so $\Pi(s)=\xi$ and $\Pi(t)=\eta$. As we have noted above, this implies $\Pi([s,t])=[\xi,\eta]$, so applying $\Pi$ to formula (3.1) we obtain
@@
The last summand in the right hand side by definition is $-\tau(\xi,\eta)$, while naturality of the fundamental derivative implies that $D_{t}s\in\Gamma(\mathcal{A}^{(0,*)}M)$ and hence $\Pi(D_{t}s)\in\Gamma(T_{\rho}M)$. From Theorem 5, we know that the first term in the right hand side is $\Pi(\nabla^{\mathcal{A}M}_{\xi}t)$ and that this is a section of $T^{i^{\prime}}_{P}M$, so (1) follows. Theorem 3.1 also tells us that, if $\eta$ represents $\sigma$, then projecting $\Pi(\nabla^{\mathcal{A}M}_{\xi}t)$ further to $\Gamma(\mathcal{V}_{i^{\prime}}M)$, we obtain $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}_{\xi}\sigma$. Since $\Pi(D_{t}s)$ lies in the kernel of the latter projection, we obtain (2). ∎
The relation to the Lie bracket is also crucial for obtaining an interpretation of the relative tractor bundles introduced in §2.3 in the case that the relative tangent bundle $T_{\rho}M$ is involutive (and additional assumptions are satisfied). As observed in §2.3, if $T_{\rho}M$ is involutive, then locally around any $x\in M$, we can find a local leaf space $\psi:U\to N$. This means that $U\subset M$ is open with $x\in U$ and $\psi$ is a surjective submersion onto a smooth manifold such that for each $y\in U$ the kernel $\ker(T_{y}\psi)\subset T_{y}M$ coincides with the fiber of the relative tangent bundle in $y$. In particular, this implies that we obtain an isomorphism $TU/T_{\rho}U\cong\psi^{*}TN$ induced by $T\psi|_{U}$. Hence there is the hope that the subbundles $T^{i^{\prime}}_{P}M\subset TM$ that contain $T_{\rho}M$ descend to subbundles in $TN$. We can nicely characterize when this happens in terms of the torsion of the Cartan geometry. Under stronger assumptions (which are satisfied in important examples), we can nicely characterize parallel sections for the relative tractor connection.
Theorem 8.
In the setting of §2.3 assume that $T_{\rho}M$ is involutive and consider a local leaf space $\psi:U\to N$ with connected fibers. Fix an index $i^{\prime}$ and consider the bundles $T^{i^{\prime}}_{P}M\subset TM$ and $T^{i^{\prime}}_{P}M/T_{\rho}M\subset TM/T_{\rho}M$.
(1) There is a smooth subbundle $T^{i^{\prime}}N\subset TN$ that corresponds to $T^{i^{\prime}}_{P}M/T_{\rho}M$ under the isomorphism $TM/T_{\rho}M\cong\psi^{*}TN$ induced by $T\psi$ if and only if the torsion $\tau$ satisfies $\tau(T_{\rho}M,T^{i^{\prime}}_{P}M)\subset T^{i^{\prime}}_{P}M$.
(2) If the condition in (1) is satisfied, then $T\psi$ induces an isomorphism $\mathcal{V}_{i^{\prime}}M|_{U}\to\psi^{*}(T^{i^{\prime}}N/T^{i^{\prime}+1}N)$. Assuming in addition that $\tau(T_{\rho}M,T^{i^{\prime}}_{P}M)\subset T^{i^{\prime}+1}_{P}M\subset T^{i^{\prime}}_{P}M$, a section of $\mathcal{V}_{i^{\prime}}M|_{U}$ is parallel for the relative tractor connection $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}$ if and only if it is the pullback of a section of $T^{i^{\prime}}N/T^{i^{\prime}+1}N$.

### v2

@@ 3.1 Restrictions of the adjoint tractor connection
A crucial feature of tractor bundles is that they carry canonical linear connections known as tractor connections. There is a uniform description of these tractor connections, based on the so-called fundamental derivative. For a geometry $(p\colon\mathcal{G}\to M,\omega)$ of type $(G,Q)$ consider any representation $\mathbb{W}$ of $Q$ and the corresponding natural vector bundle $\mathcal{W}M:=\mathcal{G}\times_{Q}\mathbb{W}$. Then smooth sections of $\mathcal{W}M$ are in bijective correspondence with smooth functions $\mathcal{G}\to\mathbb{W}$ which are $Q$-equivariant in the sense that $f(u\cdot g)=g^{-1}\cdot f(u)$ for any $g\in Q$, where in the left-hand side we use the principal right action of $Q$ on $\mathcal{G}$ and in the right-hand side the given representation on $\mathbb{W}$. In this way, sections of the adjoint tractor bundle $\mathcal{A}M$ get identified with $Q$-equivariant smooth functions $f\colon\mathcal{G}\to\mathfrak{g}$. But via the Cartan connection $\omega$, smooth functions $\mathcal{G}\to\mathfrak{g}$ are in bijective correspondence with vector fields on $\mathcal{G}$, and the equivariancy condition on $f$ is equivalent to the corresponding vector field $\xi\in\mathfrak{X}(\mathcal{G})$ being invariant under the principal right action of $Q$, i.e., $(r^{g})^{*}\xi=\xi$ for any $g\in Q$.
@@
Differentiating an equivariant function with respect to an invariant vector field produces an equivariant function, so we can view this as defining an operator $D\colon\Gamma(\mathcal{A}M)\times\Gamma(\mathcal{W}M)\to\Gamma(\mathcal{W}M)$ called the fundamental derivative, see [7, Section 1.5.8]. This has strong invariance properties and to emphasize the analogy to a covariant derivative, it is usually written as ${(s,\sigma)\!\mapsto\!D_{s}\sigma}$.
@@
While things are not formulated in this way in [10] (see the proof of Theorem 3.1 for details), relative tractor connections can be constructed in a very similar fashion. Suppose we have a bundle $\mathcal{V}M$ that is induced by a representation of $P$, where $Q\subset P\subset G$ as in Section 2.2. Then for any $g\in Q$, the subspace $\mathfrak{p}\subset\mathfrak{g}$ is invariant under $\operatorname{Ad}(g)\colon\mathfrak{g}\to\mathfrak{g}$, so it gives rise to a smooth subbundle in $\mathcal{A}M$ that is denoted by $\mathcal{A}^{\mathfrak{p}}M$ in [10]. The infinitesimal representation again induces a bundle map $\bullet\colon\mathcal{A}^{\mathfrak{p}}M\times\mathcal{V}M\to\mathcal{V}M$. As before, we conclude that the operator $\Gamma(\mathcal{A}^{\mathfrak{p}}M)\times\Gamma(\mathcal{V}M)\to\Gamma(\mathcal{V}M)$ defined by $(s,\sigma)\mapsto D_{s}\sigma+s\bullet\sigma$ depends only on $\Pi(s)$, but now $\Pi(s)$ is a section of $\mathcal{G}\times_{Q}(\mathfrak{p}/\mathfrak{q})=T_{\rho}M$. Hence, we do not obtain a full linear connection but only a partial connection $\nabla^{\rho,\mathcal{V}}\colon\Gamma\bigl{(}T_{\rho}M\bigr{)}\times\Gamma(\mathcal{V}M)\to\Gamma(\mathcal{V}M)$, which still is linear over smooth functions in the first argument and satisfies the Leibniz rule in the second argument.
In the special case of the relative tractor bundles $\mathcal{V}_{i^{\prime}}M$ constructed in Section 2.3, we prove that the relative tractor connection is induced by the adjoint tractor connection.
Theorem 3.1.
In the setting of Section 2.3, consider the subbundles $\mathcal{A}^{(i^{\prime},*)}M$. Then for ${s\in\Gamma\bigl{(}\mathcal{A}^{(0,*)}M\bigr{)}}$ and $t\in\Gamma\bigl{(}\mathcal{A}^{(i^{\prime},*)}M\bigr{)}$ with $i^{\prime}\leq 0$, we get $\nabla^{\mathcal{A}}_{\Pi(s)}t\in\Gamma\big{(}\mathcal{A}^{(i^{\prime},*)}\big{)}M$. Hence, the adjoint tractor connection restricts to an operation $\Gamma\bigl{(}T_{\rho}M\bigr{)}\times\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}\to\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$ for any $i^{\prime}<0$.
Thus, there is an induced partial connection on $\Gamma(\mathcal{V}_{i^{\prime}}M)$ and this coincides with the relative tractor connection $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}$ as obtained in [10, Section 4.3].
@@
Consider the defining formula $\nabla^{\mathcal{A}}_{\Pi(s)}t=D_{s}t+\{s,t\}$ for $s\in\Gamma\bigl{(}\mathcal{A}^{(0,*)}M\bigr{)}$ and $t\in\Gamma\smash{\bigl{(}\mathcal{A}^{(i^{\prime},*)}M\bigr{)}}$. By [7, Proposition 1.5.8], the fundamental derivative preserves natural subbundles, so the first summand in the right-hand side is a section of $\mathcal{A}^{(i^{\prime},*)}M$. On the other hand, Proposition 2.4 together with $\mathfrak{p}=\mathfrak{g}^{(0,*)}$ shows that $\bigl{[}\mathfrak{g}^{(0,*)},\mathfrak{g}^{(i^{\prime},*)}\bigr{]}\subset\mathfrak{g}^{(i^{\prime},*)}$, so the second summand of the right-hand side of the defining formula is a section of $\mathcal{A}^{(i^{\prime},*)}M$, too.
@@
Fixing $i^{\prime}<0$, we have thus obtained a partial connection on the bundle $\mathcal{A}^{(i^{\prime},*)}M$ for which the subbundle $\mathcal{A}^{(i^{\prime}+1,*)}M$ is parallel, so it descends to a well-defined partial connection on the quotient bundle $\mathcal{V}_{i^{\prime}}M=\mathcal{A}^{(i^{\prime},*)}M/\mathcal{A}^{(i^{\prime}+1,*)}M$. Thus, it remains to show that this coincides with the relative tractor connection. The construction of the latter in [10] was based on a variant of the fundamental derivative for relative natural bundles: One defines the relative adjoint tractor bundle $\mathcal{A}_{\rho}M=\mathcal{G}\times_{Q}(\mathfrak{p}/\mathfrak{p}_{+})$, so in our notation, this is just $\mathcal{A}^{(0,*)}M/\mathcal{A}^{(1,*)}M$. Given a representation $\mathbb{W}$ of $Q$ on which $P_{+}$ acts trivially, one can first restrict the fundamental derivative $D_{s}\sigma$ to $s\in\Gamma\bigl{(}\mathcal{A}^{(0,*)}M\bigr{)}$ and then observe that $D_{s}\sigma=0$ if $s\in\Gamma\bigl{(}\mathcal{A}^{(1,*)}M\bigr{)}$, to obtain a well-defined operator $D^{\rho}\colon\Gamma(\mathcal{A}_{\rho}M)\times\Gamma(\mathcal{W}M)\to\Gamma(\mathcal{W}M)$. If $\mathcal{V}M$ is a relative tractor bundle, then $\mathbb{V}$ is the restriction of a representation of $P$, so the infinitesimal representation descends to an operation $\bullet\colon\mathfrak{p}/\mathfrak{p}_{+}\times\mathbb{V}\to\mathbb{V}$. Similarly as above, the relative tractor connection is then characterized by $\nabla^{\rho,\mathcal{V}}_{\Pi(s)}\sigma=D^{\rho}_{s}\sigma+s\bullet\sigma$, where now $\Pi$ is induced by the canonical projection ${\mathfrak{p}/\mathfrak{p}_{+}\to\mathfrak{p}/\mathfrak{q}}$.
@@
But from this description the claim follows easily: Take sections $\sigma\in\Gamma(\mathcal{V}_{i^{\prime}}M)$ and $\xi\in T_{\rho}M$. For our construction, we choose representative sections $t\in\Gamma\bigl{(}\mathcal{A}^{(i^{\prime},*)}M\bigr{)}$ for $\sigma$ and $s\in\Gamma\bigl{(}\mathcal{A}^{(0,*)}M\bigr{)}$ for $\xi$ and project $D_{s}t+\{s,t\}\in\Gamma\bigl{(}\mathcal{A}^{(i^{\prime},*)}M\bigr{)}$ to the quotient bundle $\mathcal{V}_{i^{\prime}}M$. Denoting by $\underline{s}$ the projection of $s$ to $\mathcal{A}_{\rho}M$ the fact that $\mathfrak{p}_{+}$ acts trivially on $\mathfrak{g}^{(i^{\prime},*)}/\mathfrak{g}^{(i^{\prime}+1,*)}$ implies that the projection of $D_{s}t$ to $\Gamma(\mathcal{V}_{i^{\prime}}M)$ depends only on $\underline{s}$ and then by construction it coincides with $D^{\rho}_{\underline{s}}\sigma$. Similarly, since $\bigl{[}\mathfrak{g}^{(1,*)},\mathfrak{g}^{(i^{\prime},*)}\bigr{]}\subset\mathfrak{g}^{(i^{\prime}+1,*)}$ we conclude that the projection of $\{s,t\}$ to $\Gamma(\mathcal{V}_{i^{\prime}}M)$ depends only on $\underline{s}$ and coincides with $\underline{s}\bullet\sigma$. ∎
Remark 3.2.
@@
The description of a relative tractor connection as induced by the adjoint tractor connection leads to explicit descriptions in terms of the structure underlying a parabolic geometry of type $(G,Q)$. This works via the concept of Weyl structures which was introduced in [6], see also [7, Section 5]. Any parabolic geometry carries a family of Weyl structures which form an affine space modelled on the space $\Omega^{1}(M)$ of one-forms on the manifold $M$. Any Weyl structure provides a linear connection on any natural vector bundle called the Weyl connection. Once Weyl connections are understood there is a general theory providing explicit formulae for tractor connections, see [7, Section 5.1].
3.2 The relation to the Lie bracket
@@
As we have noted above, sections of the adjoint tractor bundle $\mathcal{A}M$ can be identified with $Q$-invariant vector fields on the total space $\mathcal{G}$ of the Cartan bundle. Now the subspace of $Q$-invariant vector fields is closed under the Lie bracket of vector fields, whence we obtain an induced bilinear operation on $\Gamma(\mathcal{A}M)$, which we denote by the usual symbol [,][\,,\,][ , ]. As we have noted in Section 3.1, for $s\in\Gamma(\mathcal{A}M)$, $\Pi(s)\in\mathfrak{X}(M)$ is the projection of the $Q$-invariant vector field corresponding to $s$. Hence, standard properties of the Lie bracket imply that for $s,t\in\Gamma(\mathcal{A}M)$, we get $\Pi([s,t])=[\Pi(s),\Pi(t)]$ with the Lie bracket of vector fields in the right-hand side.
@@
It is rather easy to express the Lie bracket of sections of $\mathcal{A}M$ in terms of the operations introduced above and of the curvature $\kappa$ of the Cartan connection $\omega$, see [7, Corollary 1.5.8]: For $s,t\in\Gamma(\mathcal{A}M)$, one has
$[s,t]=D_{s}t-D_{t}s-\kappa(\Pi(s),\Pi(t))+\{s,t\}.$ (3.1)
@@
Using this and Theorem 3.1, we can now give an alternative description of the relative tractor connections on the relative tractor bundles $\mathcal{V}_{i^{\prime}}M$. Recall from Section 2.3 that we have subbundles $TM\supset T^{i^{\prime}}_{P}M\supset T_{\rho}M$ such that $\mathcal{V}_{-1}M=T^{-1}_{P}M/T_{\rho}M$ and $\mathcal{V}_{i^{\prime}}M=T^{i^{\prime}}_{P}M/T^{i^{\prime}+1}_{P}M$ for ${i^{\prime}<-1}$. Hence, for a smooth section $\sigma$ of $\mathcal{V}_{i^{\prime}}M$ one can always chose a representative vector field in $\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$ that projects onto $\sigma$.
Theorem 3.3.
@@
In the setting of Section 2.3, consider vector fields $\xi\in\Gamma\bigl{(}T_{\rho}M\bigr{)}$, $\eta\in\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$ and a section $\sigma\in\Gamma(\mathcal{V}_{i^{\prime}}M)$. Let $\tau\in\Omega^{2}(M,TM)$ be the torsion of the Cartan connection $\omega$.
  • $(1)$
    The vector field $[\xi,\eta]+\tau(\xi,\eta)$ is a section of the subbundle $T^{i^{\prime}}_{P}M\subset TM$.
  • $(2)$
@@
    If $\eta$ is a representative for $\sigma$, then the projection of $[\xi,\eta]+\tau(\xi,\eta)$ to $\Gamma(\mathcal{V}_{i^{\prime}}M)$ depends only on $\sigma$ ((((and not on the choice of $\eta)$ and coincides with $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}_{\xi}\sigma$.
@@
We can choose smooth representatives $s\in\Gamma\bigl{(}\mathcal{A}^{(0,*)}M\bigr{)}$ for $\xi$ and $t\in\Gamma\bigl{(}\mathcal{A}^{(i^{\prime},*)}M\bigr{)}$ for $\eta$, so $\Pi(s)=\xi$ and $\Pi(t)=\eta$. As we have noted above, this implies $\Pi([s,t])=[\xi,\eta]$, so applying $\Pi$ to formula (3.1) we obtain
@@
The last summand in the right-hand side by definition is $-\tau(\xi,\eta)$, while naturality of the fundamental derivative implies that $D_{t}s\in\Gamma\bigl{(}\mathcal{A}^{(0,*)}M\bigr{)}$ and hence $\Pi(D_{t}s)\in\Gamma\bigl{(}T_{\rho}M\bigr{)}$. From Theorem 3.1, we know that the first term in the right-hand side is $\Pi\bigl{(}\nabla^{\mathcal{A}M}_{\xi}t\bigr{)}$ and that this is a section of $T^{i^{\prime}}_{P}M$, so (1) follows. Theorem 3.1 also tells us that, if $\eta$ represents $\sigma$, then projecting $\Pi\bigl{(}\nabla^{\mathcal{A}M}_{\xi}t\bigr{)}$ further to $\Gamma(\mathcal{V}_{i^{\prime}}M)$, we obtain $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}_{\xi}\sigma$. Since $\Pi(D_{t}s)$ lies in the kernel of the latter projection, we obtain (2). ∎
The relation to the Lie bracket is also crucial for obtaining an interpretation of the relative tractor bundles introduced in Section 2.3 in the case that the relative tangent bundle $T_{\rho}M$ is involutive (and additional assumptions are satisfied). As observed in Section 2.3, if $T_{\rho}M$ is involutive, then locally around any $x\in M$, we can find a local leaf space $\psi\colon U\to N$. This means that $U\subset M$ is open with $x\in U$ and $\psi$ is a surjective submersion onto a smooth manifold such that for each $y\in U$ the kernel $\ker(T_{y}\psi)\subset T_{y}M$ coincides with the fiber of the relative tangent bundle in $y$. In particular, this implies that we obtain an isomorphism $TU/T_{\rho}U\cong\psi^{*}TN$ induced by $T\psi|_{U}$. Hence, there is the hope that the subbundles $T^{i^{\prime}}_{P}M\subset TM$ that contain $T_{\rho}M$ descend to subbundles in $TN$. We can nicely characterize when this happens in terms of the torsion of the Cartan geometry. Under stronger assumptions (which are satisfied in important examples), we can nicely characterize parallel sections for the relative tractor connection.
Theorem 3.4.
In the setting of Section 2.3, assume that $T_{\rho}M$ is involutive and consider a local leaf space $\psi\colon U\to N$ with connected fibers. Fix an index $i^{\prime}$ and consider the bundles $T^{i^{\prime}}_{P}M\subset TM$ and $T^{i^{\prime}}_{P}M/T_{\rho}M\subset TM/T_{\rho}M$.
  1. $(1)$
    There is a smooth subbundle $T^{i^{\prime}}N\subset TN$ that corresponds to $T^{i^{\prime}}_{P}M/T_{\rho}M$ under the isomorphism $TM/T_{\rho}M\cong\psi^{*}TN$ induced by $T\psi$ if and only if the torsion $\tau$ satisfies $\tau\bigl{(}T_{\rho}M,T^{i^{\prime}}_{P}M\bigr{)}\subset T^{i^{\prime}}_{P}M$.
  2. $(2)$
    If the condition in $(1)$ is satisfied, then $T\psi$ induces an isomorphism from $\mathcal{V}_{i^{\prime}}M|_{U}$ to $\psi^{*}\bigl{(}T^{i^{\prime}}N/T^{i^{\prime}+1}N\bigr{)}$. Assuming in addition that $\tau\bigl{(}T_{\rho}M,T^{i^{\prime}}_{P}M\bigr{)}\subset T^{i^{\prime}+1}_{P}M\subset T^{i^{\prime}}_{P}M$, a section of $\mathcal{V}_{i^{\prime}}M|_{U}$ is parallel for the relative tractor connection $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}$ if and only if it is the pullback of a section of $T^{i^{\prime}}N/T^{i^{\prime}+1}N$.